If you were to walk around the perimeter of a particular building, you would cover 2476 ft. If this rectangular building is 166 ft longer than it is wide, what are its dimensions?
I know the equation P = 2L + 2W, so I think it would be something like 2476 2L + 2(W+166). At least, that's what I've tried from there. But I don't know what to do next.

Answer by jhunjiro(67)

-
Given:
P=2476
L=166+W
Solution
-
P=2L +2W
2476 = 2L + 2W
2476 = 2(166+W) + 2W
2476 = 332 + 2W + 2W
2476 -332 = 4W
2144=4W
W=536
L= 166+W
L= 166+536
L= 702
Final Answer:
L=702
W=536